# Semantic Equitable Clustering: A Simple and Effective Strategy for Clustering Vision Tokens

### v2

@@ 3.2 Semantic Equitable Clustering
Models Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
DeiT-S 22 4.6 3204 79.8 44.5 40.1 43.0
EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
SEC-DeiT-S (num_cluster=2) 22 4.3 3226 80.6 (+0.8) 47.9 (+3.4) 43.1 (+3.0) 48.0 (+5.0)
SEC-DeiT-S (num_cluster=8) 22 3.9 3528 80.1 (+0.3) 46.7 (+2.2) 42.0 (+1.9) 46.4 (+3.4)
Table 1: Comparison of EViT [35] and SEC. Inference speed are measured on the A100 GPU.
Different adaptability to vision models/MLLMs. In pruning methods like EViT, during the model’s forward pass, the number of tokens gradually decreases. Although reducing tokens does not impact classification tasks, it prevents the feature map from being restored to its original shape. This makes it difficult for EViT to be directly used with classic frameworks like SemanticFPN for downstream dense prediction tasks . In SEC, we simply group the tokens without changing their quantity, thereby preserving the integrity of the feature map. This ensures that SEC can be easily applied to downstream tasks. In Tab. 1, we use DeiT [46] as the baseline. Without introducing any other tricks, just by applying SEC to DeiT, we form SEC-DeiT and compare it with EViT-DeiT. We conducted image classification, object detection, instance segmentation and semantic segmentation based on SEC-DeiT-S. SEC not only accelerates the model but also enhances its performance. Beyond the vision tasks, we also compare the EViT with our SEC on MLLM, details can be found in the Tab. 8.
@@ 4.1 SEC for vision models
We compare our SECViT with numerous state-of-the-art models, the results are shown in Tab.4. We adopt the training strategy proposed in DeiT [46], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. SECViT consistently outperforms preceding models across all scales. Notably, SECViT-S attains a Top1-accuracy of 84.3% with a mere 27M parameters and 4.6G FLOPs. For larger models, SECViT-XL achieves a Top1-accuracy of 86.3% with 205M parameters and 36.4G FLOPs. The comparison of the models’ efficiency can be found in Appendix.
@@ Appendix A Experimental Details
ImageNet Image Classification.
We adopt the training strategy proposed in DeiT [46], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. The AdamW is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The batch-size is set to 1024, respectively. We apply the same data augmentation and regularization used in DeiT [46], including RandAugment [8] (randm9-mstd0.5-inc1) , Mixup [64] (prob = 0.8), CutMix [62] (prob = 1.0), Random Erasing (prob = 0.25). The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5/0.65 for SECViT-T/S/B/L/XL.
COCO Object Detection and Instance Segmentation.
We apply RetinaNet [36], Mask-RCNN [23], and Cascaded Mask R-CNN [2] as the detection frameworks based on the MMDetection [4]. All of our models are trained with “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings. For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS”, multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels. For both frameworks, we use the AdamW with the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. While for Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
ADE20K Semantic Segmentaion.
Based on MMSegmentation [7], we implement UperNet [55] and SemanticFPN [29] to evaluate our models’ performance on semantic segmentation. For UperNet, we follow the previous setting of Swin-Transformer [41] and train the model for 160k iterations with the input size of $512\times 512$. For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.
@@ Appendix B More experimental Results
Number of Vision Tokens Outputs by SEC.
MLLM is quite sensitive to the number of vision tokens. We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab. 14. The first row represents the speed and performance of the original LLaVA-1.5 without using SEC. Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves inference efficiency. As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
288+1 126s 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
256+1 116s 1.7×1.7$\times$1.7 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
192+1 102s 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
144+1 94s 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)

### v3

@@ 4.1 SEC for vision models
We compare our SECViT with numerous state-of-the-art models, the results are shown in Tab.3. We adopt the training strategy proposed in DeiT [43], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. SECViT consistently outperforms preceding models across all scales. Notably, SECViT-S attains a Top1-accuracy of 84.3% with a mere 27M parameters and 4.6G FLOPs. The comparison of the models’ efficiency can be found in Appendix.
@@ 4.3 Ablation Study
4.4 Visualization of SEC
Figure 5: Visualization for SEC.
Number of Vision Tokens Outputs by SEC.
MLLM is quite sensitive to the number of vision tokens. We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab. 11. The first row represents the speed and performance of the original LLaVA-1.5 without using SEC. Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves training efficiency. As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
576+1 21h 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
288+1 14h 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
256+1 13h 1.6×1.6$\times$1.6 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
192+1 11h 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
144+1 10h 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)
@@ 5 Conclusion
6 Acknowledgements
This work is partially funded by Beijing Natural Science Foundation (4252054), Youth Innovation Promotion Association CAS(Grant No.2022132), Beijing Nova Program(20230484276), and CCF-Kuaishou Large Model Explorer Fund (NO. CCF-KuaiShou 2024005).
[…]
Supplementary Material